Perpetual Inventory Using LIFO

Beginning inventory, purchases, and sales for Item 88-HX are as follows:

Jul. 1 Inventory 86 units @ $15
8 Sale 69 units
15 Purchase 95 units @ $17
27 Sale 80 units

Assuming a perpetual inventory system and using the last-in, first-out (LIFO) method, determine (a) the cost of goods sold on Jul. 27 and (b) the inventory on Jul. 31.

a. Cost of goods sold on Jul. 27 $
b. Inventory on Jul. 31 $
Answer #1

As per LIFO cost of goods sold:

Cost of goods sold= (69*15)+(80*17)

Cost of goods sold= 1035+1360= 2395

Cost of goods sold= 2395

Calculation of ending inventory:

Ending inventory= (17*15)+(15*17)= 510

Ending inventory is 510